# DiffGAN: A Test Generation Approach for Differential Testing of Deep Neural Networks for Image Analysis

### v3

@@ 5 Empirical Results
We also manually analyzed the types of triggering inputs generated by DiffGAN and DRFuzz by randomly sampling 1,000 labeled valid triggering images for each method and model pair. Each input was categorized into one of the following types: (1) only Model A produces the correct label, (2) only Model B produces the correct label, and (3) both models produce incorrect but different labels. Here, Model A refers to the first model in each pair, and Model B to the second. This categorization enabled us to quantify the nature of disagreements revealed by the generated inputs. As shown in Table XII, both DiffGAN and DRFuzz exhibit all disagreement types, however, in different proportions. Indeed, DiffGAN generates significantly fewer inputs where both models are incorrect. This is practically important as we want to generate as many inputs that distinguish the two models’ behavior as possible, regarding conditions when they can be trusted, in order to learn which model to trust for a given input. In turn, this is useful to devise accurate model selectors, for example, at run-time. In conclusion, these results further explain why DiffGAN yields better model selectors than DRFuzz.
@@
To compare the accuracy of the RF-based model selection systems, we categorized the images into three groups based on whether one model predicted correctly or if both models failed. We excluded images where both models failed from further analysis due to their low occurrence and lack of relevance to the comparison (i.e., if both models fail, it does not matter which classification model is selected). We trained the RF classifiers using features extracted from a pre-trained VGG16 model, rather than raw pixel data, in an attempt to improve accuracy. The VGG16 model provides feature vectors that capture more abstract and informative patterns than pixel-based features, which is expected to improve the RF model’s ability to distinguish between correct and incorrect predictions. For each method and model pair, we selected 700 valid labeled images and used their VGG16 features to train eight RF-based selection models. To ensure a fair comparison between the RF models trained on different datasets in each scenario, we need a common test dataset. For each scenario, we selected 100 images from the DiffGAN-generated test set and 100 from the DRfuzz-generated test set, creating a combined test set of 200 images. This test set allowed us to evaluate the performance of the RF models using the same test set in an unbiased manner. Additionally, we applied the same feature extraction process as for the training sets.
@@
We report the accuracy levels of each of the four selection model we trained in Figure 5. As clearly visible, the RF-based selection models trained using triggering inputs generated with DiffGAN are far more accurate than those trained with inputs from DRfuzz. For instance, in the first scenario involving two RF-based selection models built using classifiers Model 1 and Model 2 from Table I, the RF model trained with DRfuzz achieved an accuracy of 78%, whereas the one trained with DiffGAN reached a much higher accuracy of 87%. Overall, the accuracy levels for the RF models trained with DiffGAN range between 79% and 87% compared to a range of 70% to 78% for those with DRfuzz inputs. This highlights that DiffGAN produces more effective triggering inputs for training, allowing the RF model to better capture the conditions under which one classifier outperforms the other. As a result, RF models trained using DiffGAN inputs are more accurate in selecting the best-performing classifier, leading to a better model selection mechanism.
@@
This improved accuracy can be attributed to several factors. DiffGAN leverages the combination of GAN and NSGA-II to generate highly diverse triggering inputs. Unlike DRfuzz, which applies mutation rules to existing seed inputs, DiffGAN explores a significantly broader latent input space, enabling the generation of more diverse inputs that expose more pronounced behavioral differences between the models. These behavioral differences are crucial for the RF model to learn the conditions when each classifier performs better. Additionally, DiffGAN optimizes its inputs to both maximize diversity and reveal discrepancies, which also enables the RF model to develop a more accurate understanding of the conditions under which one classifier outperforms the other. Though the accuracy of RF models can still be improved, our main goal here was to show the usefulness of DiffGAN for one important application, not to optimize model selection. The latter would require the labeling of a much larger dataset and possibly the fine-tuning of the RF models or experimenting with other machine learning techniques, which is out of the scope of this paper.
Answer to RQ4: The ML-based model selection mechanisms trained with DiffGAN triggering inputs are more accurate in selecting the best-performing classifier among alternative models, leading to better model selection mechanisms. This highlights that DiffGAN produces more triggering inputs that are more effective for training, allowing the RF model to better capture the conditions under which one classifier outperforms the other. In turn, this can lead to more accurate voting mechanisms at inference time and thus higher classification accuracy.
@@ 6 Discussions
Relevance of our Approach: We present in this paper DiffGAN, a test generation approach for the differential testing of DNN models for image classification. DiffGAN leverages NSGA-II, a multi-objective optimization algorithm, to effectively explore the latent space of a GAN and generate diverse and valid triggering images. Unlike traditional mutation-based approaches that rely on modifying seed inputs [8, 9, 10], DiffGAN directly explores the high-dimensional GAN latent space, allowing it to generate entirely new images with novel features that can trigger behavioral disagreements between the models under test. NSGA-II plays a crucial role by balancing two fitness functions: (1) maximizing divergence—ensuring, to generate inputs leading to significant differences in models outputs, and (2) maximizing diversity, thus reducing the generation of redundant images. Our ablation study shows that these two fitness functions are complementary and both essential for the generation of triggering images. By integrating GAN-based test inputs generation with NSGA-II-driven optimization, DiffGAN not only generates more triggering inputs than the SOTA approach but also ensures that these inputs are valid and more diverse.
@@
Practical Applicability: The black-box nature of DiffGAN makes it agnostic to the models under test, as it does not require any assumptions on their architectural similarity. This flexibility allows DiffGAN to be applied across diverse practical scenarios where models with similar accuracy levels are developed independently or trained using different data and optimization techniques. The ability of DiffGAN to generate diverse and valid triggering inputs makes it a practical approach for evaluating and comparing DNN models with similar accuracy levels. DiffGAN is therefore particularly valuable for model selection, regression testing, and model compression, which are common application scenarios in ML development pipelines. In model selection, the triggering inputs generated by DiffGAN can help identify behavioral discrepancies enabling fine-grained analysis beyond accuracy metrics, ensuring the selection of the most adequate model. In regression testing, DiffGAN can be used to generate cases where an updated model underperforms compared to its predecessor. In model compression validation, DiffGAN can be used to understand when compressed models fare less well than their original versions and whether this is acceptable. Another important application of DiffGAN is ML-based model selection, a scenario informed by our industrial partner. In some autonomous vehicles, classifiers with similar accuracy levels are dynamically selected on the fly to classify sensor-captured images. Since different classifiers specialize in distinct regions of the feature space, selecting the most suitable one for each input improves reliability. Additionally, certain classifiers require more computational resources than others, making dynamic selection crucial for optimizing inference time and energy consumption, particularly in embedded systems with limited resources. By generating diverse inputs that expose model discrepancies, DiffGAN enhances the accuracy of model selection mechanisms since their training dataset is enriched by such inputs and is more informative. The relevance of this application is further supported by existing literature on ML-based model selection [51, 52].
@@
Furthermore, our evaluation includes both highly accurate model pairs (with accuracy above 98%) and relatively less accurate ones (ranging from 80% to 97%) to enhance the generalizability of our results. For highly accurate models, the generation of additional triggering inputs is considerably more challenging due to the scarcity of disagreement regions in the input space. For less accurate models, while behavioral disagreements may be more frequent, the key challenge remains in generating inputs that are not only valid but also diverse, so as to avoid trivial or uninformative disagreements. As shown in our ablation study (Table III), maximizing only divergence without accounting for diversity leads to large numbers of triggering inputs, but many are similar, offering limited insight into the boundary behaviors of the models. While the original test dataset contains triggering inputs for each model pair, these inputs could be redundant and typically do not sufficiently capture the diverse conditions under which behavioral differences between models occur. Relying solely on these inputs would not provide a comprehensive view of how and when models diverge in their predictions. DiffGAN addresses these challenges by generating, on average, four times more triggering inputs than the SOTA baseline, while also significantly improving both the validity and the diversity of these inputs. This richer and more diverse set of valid triggering inputs enables a more comprehensive behavioral analysis of model pairs.
@@
As shown in Figure 6, we found that DiffGAN scales linearly with increasing dataset size, with a training time ranging from 371 to 607 seconds. We note that this increase remains practically acceptable, as GAN training is a one-time process per dataset, and differential testing is neither frequent nor a real-time task. Thus, training costs do not significantly impact the feasibility of using DiffGAN for large-scale differential testing. Additionally, it is important to clarify that DiffGAN does not strictly require training a GAN from scratch, as it treats GANs (or any other generative model) as black-box components. This means that, for large-scale datasets such as ImageNet, DiffGAN can leverage pretrained GANs or generators that have already been trained on such datasets [74, 75]. This would reduce the computational burden of training new GANs while maintaining the ability of DiffGAN to generate diverse and valid test cases. We consider incorporating pretrained GANs and evaluating their effect on triggering input quality and computational savings as future work.
@@ 7 Threats to Validity
In this section, we discuss the threats to the validity of our study and describe how we mitigated them.
@@ 8 Related Work
Tian et al. [9] proposed Dflare, a search-based test generation approach that iteratively applies a series of mutation operations to a given seed image until a triggering input is found. They rely on Markov Chains and divergence-based fitness function to guide the selection of mutation operators [9]. The fitness function is designed to prioritize mutated inputs that (1) maximize differences between the output probability vectors of two models or (2) trigger new, previously unobserved probability vectors in the models under test.
@@
The above test generation approaches heavily rely on mutation operations applied on seed images, thus limiting the process to mutated versions of existing inputs. This dependence significantly narrows the range of test scenarios, as it prevents the introduction of entirely new features (i.e, features that are not present in the initial testing dataset) into the testing process. In contrast, DiffGAN enhances its differential testing effectiveness by generating diverse and new test input scenarios.
@@
Diverget [6] and DeepEvolution [7] are two white-box differential testing approaches for quantized DNN models. They leverage multi-objective search algorithms and domain-specific mutation operators to generate triggering inputs. The generation is guided by two fitness functions. The former fitness function maximizes the divergence in output probability vectors between the models under test, while the latter aims to maximize the coverage of neuron activations within those models. While DeepEvolution is designed to test both original and quantized DNN models for image classification tasks, Diverget specifically targets DNN models used for hyperspectral image classification [6].

### v4

@@ 5 Empirical Results
We also manually analyzed the types of triggering inputs generated by DiffGAN and DRfuzz by randomly sampling 1,000 labeled valid triggering images for each method and model pair. Each input was categorized into one of the following types: (1) only Model A produces the correct label, (2) only Model B produces the correct label, and (3) both models produce incorrect but different labels. Here, Model A refers to the first model in each pair, and Model B to the second. This categorization enabled us to quantify the nature of disagreements revealed by the generated inputs. As shown in Table XII, both DiffGAN and DRfuzz exhibit all disagreement types, however, in different proportions. Indeed, DiffGAN generates significantly fewer inputs where both models are incorrect. This is practically important as we want to generate as many inputs that distinguish the two models’ behavior as possible, regarding conditions when they can be trusted, to learn which model to trust for a given input. In turn, this is useful to devise accurate model selectors, for example, at run-time. In conclusion, these results further explain why DiffGAN yields better model selectors than DRfuzz.
@@
To compare the accuracy of the RF-based model selection systems, we categorized the images into three groups based on whether one model predicted correctly or if both models failed. We excluded images where both models failed from further analysis due to their low occurrence and lack of relevance to the comparison (i.e., if both models fail, it does not matter which classification model is selected). We trained the RF classifiers using features extracted from a pre-trained VGG16 model, rather than raw pixel data, in an attempt to improve accuracy. The VGG16 model provides feature vectors that capture more abstract and informative patterns than pixel-based features, which is expected to improve the RF model’s ability to distinguish between correct and incorrect predictions. For each method and model pair, we selected 700 valid labeled images and used their VGG16 features to train eight RF-based selection models. To ensure a fair comparison between the RF models trained on different datasets in each scenario, we need a common test dataset. For each scenario, we selected 100 images from the DiffGAN-generated test set and 100 images from the DRfuzz-generated test set, resulting in a combined test set of 200 images. This test set allowed us to evaluate the performance of the RF models using the same test set in an unbiased manner. Additionally, we applied the same feature extraction process as for the training sets.
@@
We report the accuracy levels of each of the four selection model we trained in Figure 5. As clearly visible, the RF-based selection models trained using triggering inputs generated with DiffGAN are far more accurate than those trained with inputs from DRfuzz. For instance, in the first scenario involving two RF-based selection models built using classifiers Model 1 and Model 2 from Table I, the RF model trained with DRfuzz achieved an accuracy of 78%, whereas the one trained with DiffGAN reached a much higher accuracy of 87%. Overall, the accuracy levels for the RF models trained with DiffGAN range between 79% and 87% compared to a range of 70% to 78% for those with DRfuzz inputs. This highlights that DiffGAN produces more effective triggering inputs for training, allowing the RF model to capture better the conditions under which one classifier outperforms the other. As a result, RF models trained using DiffGAN inputs are more accurate in selecting the best-performing classifier, resulting in a more effective model selection mechanism.
@@
This improved accuracy can be attributed to several factors. DiffGAN leverages the combination of GAN and NSGA-II to generate highly diverse triggering inputs. Unlike DRfuzz, which applies mutation rules to existing seed inputs, DiffGAN explores a significantly broader latent input space, enabling the generation of more diverse inputs that expose more pronounced behavioral differences between the models. These behavioral differences are crucial for the RF model to learn the conditions when each classifier performs better. Additionally, DiffGAN optimizes its inputs to both maximize diversity and reveal discrepancies, which also enables the RF model to develop a more accurate understanding of the conditions under which one classifier outperforms the other. Though the accuracy of RF models can still be improved, our main goal here was to show the usefulness of DiffGAN for one important application, not to optimize model selection. The latter would require labeling a much larger dataset and possibly fine-tuning the RF models or experimenting with other machine learning techniques, which is outside the scope of this paper.
Answer to RQ4: The ML-based model selection mechanisms trained with DiffGAN triggering inputs are more accurate in selecting the best-performing classifier among alternative models, leading to better model selection mechanisms. This highlights that DiffGAN produces more triggering inputs that are more effective for training, allowing the RF model to capture better the conditions under which one classifier outperforms the other. In turn, this can lead to more accurate voting mechanisms at inference time and thus higher classification accuracy.
@@ 6 Discussions
Relevance of our Approach: We present in this paper DiffGAN, a test generation approach for the differential testing of DNN models for image classification. DiffGAN leverages NSGA-II, a multi-objective optimization algorithm, to effectively explore the latent space of a GAN and generate diverse and valid triggering images. Unlike traditional mutation-based approaches that rely on modifying seed inputs [8, 9, 10], DiffGAN directly explores the high-dimensional GAN latent space, allowing it to generate entirely new images with novel features that can trigger behavioral disagreements between the models under test. NSGA-II plays a crucial role by balancing two fitness functions: (1) maximizing divergence—ensuring, to generate inputs leading to significant differences in models outputs, and (2) maximizing diversity, thus reducing the generation of redundant images. Our ablation study shows that these two fitness functions are complementary and both essential for generating triggering images. By integrating GAN-based test inputs generation with NSGA-II-driven optimization, DiffGAN not only generates more triggering inputs than the SOTA approach but also ensures that these inputs are valid and more diverse.
@@
Practical Applicability: The black-box nature of DiffGAN makes it agnostic to the models under test, as it does not require any assumptions on their architectural similarity. This flexibility allows DiffGAN to be applied across diverse practical scenarios where models with similar accuracy levels are developed independently or trained using different data and optimization techniques. The ability of DiffGAN to generate diverse and valid triggering inputs makes it a practical approach for evaluating and comparing DNN models with similar accuracy levels. DiffGAN is therefore particularly valuable for model selection, regression testing, and model compression, which are common application scenarios in ML development pipelines. In model selection, the triggering inputs generated by DiffGAN can help identify behavioral discrepancies, enabling fine-grained analysis beyond accuracy metrics, ensuring the selection of the most adequate model. In regression testing, DiffGAN can be used to generate cases where an updated model underperforms compared to its predecessor. In model compression validation, DiffGAN can be used to understand when compressed models fare less well than their original versions and whether this is acceptable. Another important application of DiffGAN is ML-based model selection, a scenario informed by our industrial partner. In some autonomous vehicles, classifiers with similar accuracy levels are dynamically selected on the fly to classify sensor-captured images. Since different classifiers specialize in distinct regions of the feature space, selecting the most suitable one for each input improves reliability. Additionally, certain classifiers require more computational resources than others, making dynamic selection crucial for optimizing inference time and energy consumption, particularly in embedded systems with limited resources. By generating diverse inputs that expose model discrepancies, DiffGAN enhances the accuracy of model selection mechanisms since their training dataset is enriched by such inputs and is more informative. The relevance of this application is further supported by existing literature on ML-based model selection [51, 52].
@@
Furthermore, our evaluation includes both highly accurate model pairs (with accuracy above 98%) and relatively less accurate ones (ranging from 80% to 97%) to enhance the generalizability of our results. For highly accurate models, the generation of additional triggering inputs is considerably more challenging due to the scarcity of disagreement regions in the input space. For less accurate models, while behavioral disagreements may be more frequent, the key challenge remains in generating inputs that are not only valid but also diverse, to avoid trivial or uninformative disagreements. As shown in our ablation study (Table III), maximizing only divergence without accounting for diversity leads to large numbers of triggering inputs, but many are similar, offering limited insight into the boundary behaviors of the models. While the original test dataset contains triggering inputs for each model pair, these inputs could be redundant and typically do not sufficiently capture the diverse conditions under which behavioral differences between models occur. Relying solely on these inputs would not provide a comprehensive view of how and when models diverge in their predictions. DiffGAN addresses these challenges by generating, on average, four times more triggering inputs than the SOTA baseline, while also significantly improving both the validity and the diversity of these inputs. This richer and more diverse set of valid triggering inputs enables a more comprehensive behavioral analysis of model pairs.
@@
As shown in Figure 6, we found that DiffGAN scales linearly with increasing dataset size, with a training time ranging from 371 to 607 seconds. We note that this increase remains practically acceptable, as GAN training is a one-time process per dataset, and differential testing is neither frequent nor a real-time task. Thus, training costs do not significantly impact the feasibility of using DiffGAN for large-scale differential testing.
Additionally, it is important to clarify that DiffGAN does not strictly require training a GAN from scratch, as it treats GANs (or any other generative model) as black-box components. This means that, for large-scale datasets such as ImageNet, DiffGAN can leverage pretrained GANs or generators that have already been trained on such datasets [74, 75]. This would reduce the computational burden of training new GANs while maintaining the ability of DiffGAN to generate diverse and valid test cases. We consider incorporating pretrained GANs and evaluating their effect on triggering input quality and computational savings as future work.
@@ 7 Threats to Validity
In this section, we discuss the potential threats to the validity of our study and describe the measures we took to mitigate them.
@@ 8 Related Work
Tian et al. [9] proposed DFlare, a search-based test generation approach that iteratively applies a series of mutation operations to a given seed image until a triggering input is found. They rely on Markov Chains and a divergence-based fitness function to guide the selection of mutation operators [9]. The fitness function is designed to prioritize mutated inputs that (1) maximize differences between the output probability vectors of two models or (2) trigger new, previously unobserved probability vectors in the models under test.
@@
The above test generation approaches heavily rely on mutation operations applied to seed images, thus limiting the process to mutated versions of existing inputs. This dependence significantly narrows the range of test scenarios, as it prevents the introduction of entirely new features (i.e, features that are not present in the initial testing dataset) into the testing process. In contrast, DiffGAN enhances its differential testing effectiveness by generating diverse and new test input scenarios.
@@
Diverget [6] and DeepEvolution [7] are two white-box differential testing approaches for quantized DNN models. They leverage multi-objective search algorithms and domain-specific mutation operators to generate triggering inputs. Two fitness functions guide the generation. The former fitness function maximizes the divergence in output probability vectors between the models under test, while the latter aims to maximize the coverage of neuron activations within those models. While DeepEvolution is designed to test both original and quantized DNN models for image classification tasks, Diverget specifically targets DNN models used for hyperspectral image classification [6].